OpenStudy (anonymous):

Two trains leave stations 546 miles apart at the same time and travel toward each other. One train travels at 95 miles per hour while the other travels at 115 miles per hour. How long will it take for the two trains to meet? Do not do any rounding.

OpenStudy (anonymous):

95+115x=546

ganeshie8 (ganeshie8):

you're almost there!

OpenStudy (anonymous):

at least im close

ganeshie8 (ganeshie8):

speed * time = distance

OpenStudy (anonymous):

95*115*546=x

ganeshie8 (ganeshie8):

In `x` hours the first train travels a distance of `95x` In `x` hours the second train travels a distance of `115x` when they meet they both would have travelled a distance of 546 miles : \[95x + 115x = 546\]

ganeshie8 (ganeshie8):

solve \(x\)

OpenStudy (anonymous):

x=2.6

ganeshie8 (ganeshie8):

yes! so it takes 2.6 hours for them to collide